The perimeter of a rectangle is 142 m.The length is 2 m more than twice the width.Find the dimensions. What is the length and width?
Answer by palanisamy(496)

Let the width of the rectangle = x
Given, the length is 2 m more than twice the width.
So the length = 2x+2
Perimeter is 2(2x+2+x) = 142
(3x+2) = 142/2
3x+2 = 71
3x = 71-2
3x = 69
x = 69/3
x = 23m
Therefore width = 23 m
length= 2x+2 = 2*23+2 = 48m